Prove the identity:
cos^2 y - sin^2 y = 1-2sin^2 y

Answer by ikleyn(52790)

.
Prove the identity:
cos^2(y) - sin^2(y) = 1 - 2sin^2(y).
~~~~~~~~~~~~~~~~~

  cos^2(y) - sin^2(y) =      (substitute here cos^2(y) = 1 - sin^2(y) to get) 


= 1 - 2*sin^2(y).

QED.